OpenStudy (anonymous):

Find the missing coordinate of , using the fact that lies on the unit circle in the given quadrant. P(?,1/13) quadrant 11 P(2/3,?) quadrant 1

OpenStudy (anonymous):

like that, see attachment

OpenStudy (dumbcow):

use pythagorean theorem for unit circle radius always equals 1 x^2 + (1/13)^2 = 1 solve for x the quadrant will tell you whether its pos or neg

OpenStudy (anonymous):

so what would be the first one?

OpenStudy (dumbcow):

x^2 = 1 - (1/13)^2 solve for x quad 2 means x is neg 

OpenStudy (anonymous):

for x i got approximatly .9970 but how would i put that in fraction form?

OpenStudy (dumbcow):

1/13 ^2 = 1/169 1 = 169/169 --> 169/169 - 1/169 = 168/169 x^2 = 168/169 x = sqrt168/13 = 2sqrt42/13

OpenStudy (anonymous):

and it will be negative bec its the 2nd quadrant right?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

and what about the second one?

OpenStudy (anonymous):

x = sqrt ( 1 - 1/169) = - sqrt (168)/ 13 =>( - sqrt (168)/ 13 , 1/13) y = sqrt ( 1 - 4/9) = sqrt (5)/3 => ( 2/3, sqrt (5)/3)